# An effective control of large systems of active particles: An application to evacuation problem

### v2

@@ 4 Experimental results
In our simulations we observe that the rescuer-leader firstly learned to walk around the walls, then it learned to collect and escort individuals to the exit; finally it learned to do this as fast as possible. In Figure 4 we show the learning kinetics – the evolution with time of the average reward of a training episode [^4] The averaging over time makes the results more interpretable and visually clear due to smoothing short-term fluctuations. (we use here an exponential moving average [67]). As it follows from the figure, the learning time is about $6\cdot 10^{4}$ time-steps – after that training time the leader ”knows” how to evacuate people in the most fast and effective way. In Figure 4 we also show the learning curve for RL with basic feedforward architecture (without “pseudo-gravitational” encoding), as well as for one of the most popular, state-of-the-art, transformer-based architecture [65]. As it is clearly seen from Figure 4 (as well from its counterpart, Figure 6 in the Appendix), the new algorithm with the “pseudo-gravitational” encoding substantially outperforms the algorithm with the transformer architecture. Moreover, it dramatically outperforms the RL with basic feedforward architecture, which essentially fails to perform an effective evacuation. In addition to higher evacuation efficiency, our method also exhibits faster and more stable convergence compared to the baselines. While transformer-based models require substantially more training steps and feed-forward models fail to converge, the pseudo-gravitational encoding consistently reaches effective strategies (see Figure 6). Such advantages of the new algorithm have been demonstrated for different scenarios – for 60 and 15 individuals in a hazardous zone, which is demonstrated in Figure 4 and Figure 6. Interestingly, our experiments show that $\alpha=3$ is the optimal value for the parameter $\alpha$ of the “pseudo-gravitational” potential (see Appendix C and Figure 7).
@@ 5 Discussion and conclusion
We believe that the integration of our “pseudo-gravitational” encoder into the neural architecture search (NAS) frameworks [68, 69] (as well as ZeroNAS [70], and DS-Net++ [71]) is a promising direction for the future studies of an effective evacuation. Furthermore, the Umbrella RL [72] also has great potential in this context. Finally, the multi-agent version of the evacuation problem, with the MARLib [73] framework, can be considered – it would provide a powerful tool by incorporating a wide range of multi-agent RL algorithms and ANN architectures. Utilizing NAS frameworks, however, would be computationally very costly, which remains a serious challenge for future studies.
@@ Acknowledgments
The work was supported by the Russian Science Foundation grant № 25-71-30008, https://rscf.ru/project/25-71-30008/. The authors also acknowledge the usage of the supercomputer Zhores [74].
@@
Application of RL to active matter seems to be very fruitful, see e.g. [9]. Here we present definitions and brief outline of the RL, applied for the generalized Vicsek model, that is, for an effective leader-guided evacuation of people from a hazardous place.
@@
The action policy $\pi(a|s)$ is optimized via stochastic gradient ascent, by a policy gradient algorithm, see e.g. [16]. Below we briefly sketch this approach, illustrating it for the Actor-Critic algorithm, which is simpler to follow. We assume, that the action policy corresponds to an ANN, specified by the learnable parameters $\theta$. Then, a sequence of episodes of the system evolution is simulated and the policy parameters are updated as
@@
is the stochastic estimate of the gradient of the expected return, $J$, defined by Equation 12. In the above equation $A$ is the advantage function estimate [75]. It is calculated from the reward received during an episode simulation. The advantage function serves to damp the gradient variance in the stochastic policy gradient estimate. This function, in its simplest implementation, reads,
@@
The first term in the right-hand side (r.h.s.) in Equation 15 is obtained in direct simulations of the environment with a current learnable policy $\pi$. From the definition of the value function $V(s)$, e.g. [16], it follows, that it may be estimated as the minimum of the expression,
@@
The gradient is calculated using automatic differentiation packages; for more detail see the original paper [76], where A2C is discussed. In practice, however, we utilize a more powerful RPO algorithm [63] that is a robust version of PPO [77] algorithm. The main idea of PPO coincides with that of A2C, but it contains some parts, which significantly improve the speed and convergence. Currently, the PPO is a widely spread, state-of-the-art RL algorithm. Schematically, the training with pseudo-gravitational encoding, is present in the Algorithm 2.
@@ Appendix B Average angle/direction
The dynamic rules for the Vicsek model require computation of the average direction of the moving neighbouring individuals. Here one needs to apply correct summing rules for the angles. For instance, for two particles, the expression for the average angle reads [53]:
@@
and corresponds to the standard formulation of the Vicsek model [5]. We use the angle notation here for compactness only. It is important to stress that a simple arithmetic mean of angles is not well-defined, since its result depends on the chosen branch of the angle interval and may yield contradictory outcomes. By contrast, averaging in the vector form and then computing the argument ensures a unique definition of the mean direction. In degenerate symmetric cases (when the average vector has zero magnitude), no prevailing orientation exists and the new direction is undefined which is also reflected in the equation.

### v3

@@ 4 Experimental results
In our simulations we observe that the rescuer-leader firstly learned to walk around the walls, then it learned to collect and escort individuals to the exit; finally it learned to do this as fast as possible. In Figure 4 we show the learning kinetics – the evolution with time of the average reward of a training episode [^4] The averaging over time makes the results more interpretable and visually clear due to smoothing short-term fluctuations. (we use here an exponential moving average [brown1956exponential]). As it follows from the figure, the learning time is about $6\cdot 10^{4}$ time-steps – after that training time the leader ”knows” how to evacuate people in the most fast and effective way. In Figure 4 we also show the learning curve for RL with basic feedforward architecture (without “pseudo-gravitational” encoding), as well as for one of the most popular, state-of-the-art, transformer-based architecture [vaswani2017attention]. As it is clearly seen from Figure 4 (as well from its counterpart, Figure 6 in the Appendix), the new algorithm with the “pseudo-gravitational” encoding substantially outperforms the algorithm with the transformer architecture. Moreover, it dramatically outperforms the RL with basic feedforward architecture, which essentially fails to perform an effective evacuation. In addition to higher evacuation efficiency, our method also exhibits faster and more stable convergence compared to the baselines. While transformer-based models require substantially more training steps and feed-forward models fail to converge, the pseudo-gravitational encoding consistently reaches effective strategies (see Figure 6). Such advantages of the new algorithm have been demonstrated for different scenarios – for 60 and 15 individuals in a hazardous zone, which is demonstrated in Figure 4 and Figure 6. Interestingly, our experiments show that $\alpha=3$ is the optimal value for the parameter $\alpha$ of the “pseudo-gravitational” potential (see Appendix C and Figure 7).
@@ 5 Discussion and conclusion
We believe that the integration of our “pseudo-gravitational” encoder into the neural architecture search (NAS) frameworks [ren2021comprehensive, cheng2020hierarchical] (as well as ZeroNAS [yan2021zeronas], and DS-Net++ [li2022ds]) is a promising direction for the future studies of an effective evacuation. Furthermore, the Umbrella RL [Nuzhin_2024_CNSNS] also has great potential in this context. Finally, the multi-agent version of the evacuation problem, with the MARLib [hu2023marllib] framework, can be considered – it would provide a powerful tool by incorporating a wide range of multi-agent RL algorithms and ANN architectures. Utilizing NAS frameworks, however, would be computationally very costly, which remains a serious challenge for future studies.
@@ Acknowledgments
The work was supported by the Russian Science Foundation grant № 25-71-30008, https://rscf.ru/project/25-71-30008/. The authors also acknowledge the usage of the supercomputer Zhores [zhoresSuper].
@@
Application of RL to active matter seems to be very fruitful, see e.g. [Egor2021]. Here we present definitions and brief outline of the RL, applied for the generalized Vicsek model, that is, for an effective leader-guided evacuation of people from a hazardous place.
@@
The action policy $\pi(a|s)$ is optimized via stochastic gradient ascent, by a policy gradient algorithm, see e.g. [RL2]. Below we briefly sketch this approach, illustrating it for the Actor-Critic algorithm, which is simpler to follow. We assume, that the action policy corresponds to an ANN, specified by the learnable parameters $\theta$. Then, a sequence of episodes of the system evolution is simulated and the policy parameters are updated as
@@
is the stochastic estimate of the gradient of the expected return, $J$, defined by Equation 12. In the above equation $A$ is the advantage function estimate [schulman2015high]. It is calculated from the reward received during an episode simulation. The advantage function serves to damp the gradient variance in the stochastic policy gradient estimate. This function, in its simplest implementation, reads,
@@
The first term in the right-hand side (r.h.s.) in Equation 15 is obtained in direct simulations of the environment with a current learnable policy $\pi$. From the definition of the value function $V(s)$, e.g. [RL2], it follows, that it may be estimated as the minimum of the expression,
@@
The gradient is calculated using automatic differentiation packages; for more detail see the original paper [mnih2016asynchronous], where A2C is discussed. In practice, however, we utilize a more powerful RPO algorithm [rahman2022robust] that is a robust version of PPO [schulman2017proximal] algorithm. The main idea of PPO coincides with that of A2C, but it contains some parts, which significantly improve the speed and convergence. Currently, the PPO is a widely spread, state-of-the-art RL algorithm. Schematically, the training with pseudo-gravitational encoding, is present in the Algorithm 2.
@@ Appendix B Average angle/direction
The dynamic rules for the Vicsek model require computation of the average direction of the moving neighbouring individuals. Here one needs to apply correct summing rules for the angles. For instance, for two particles, the expression for the average angle reads [labasken]:
@@
and corresponds to the standard formulation of the Vicsek model [VicsekINITIAL]. We use the angle notation here for compactness only. It is important to stress that a simple arithmetic mean of angles is not well-defined, since its result depends on the chosen branch of the angle interval and may yield contradictory outcomes. By contrast, averaging in the vector form and then computing the argument ensures a unique definition of the mean direction. In degenerate symmetric cases (when the average vector has zero magnitude), no prevailing orientation exists and the new direction is undefined which is also reflected in the equation.